OpenStudy (anonymous):

I really need help! Given the displacement function s(t)=2sint, determine the velocity and acceleration functions.

OpenStudy (richyw):

velocity is the time derivative of position. acceleration is the time derivative of velocity.\[s(t)=2\sin t\]\[v(t)=s'(t)=2\cos t\]\[a(t)=v'(t)=-2\sin t\]